Help!!! I'm having a very hard time with this exam review.
Determine the linear function whose graph is a line with a slope of 5/7 and contains the point (-1,5). Thanks!

Answer by ankor@dixie-net.com(22740)

Determine the linear function whose graph is a line with a slope of 5/7 and contains the point (-1,5).
:
There is a formula called the "Point/slope" formula. Memorize this and you won't have any problem with it: y - y1 = m(x - x1)
:
In your problem the slope (m) = 5/7; x1 = -1; y1 = 5; substitute these values.
:
y - 5 = (5/7)(x - (-1))
y - 5 = (5/7)(x + 1); remember minus a minus is a plus
y = 5 = (5/7)x + (5/7); multiply what's inside the brackets
y = (5/7)x + 5/7 + 5
y = (5/7)x + 5/7 + 35/7; changed 5 to 35/7
y = (5/7)x + 40/7
:
If they want in the standard form, multiply the equation by 7 and rearrange:
-5x + 7y = 40
:
Remember you can always check to see if you are right by substituting -1
for x and 5 for y in the above equation to see if it = 40
:
Did this help?